13. What is the molar solubility of nickel hydroxide, Ni(OH)2? Ksp = 6.5 x 10-18
Answer #1

At equilibrium:

Ni(OH)2 <----> Ni2+ + 2 OH-

   s 2s

Ksp = [Ni2+][OH-]^2

6.5*10^-18=(s)*(2s)^2

6.5*10^-18= 4(s)^3

s = 1.176*10^-6 M

Answer: 1.2*10^-6 M